OpenStudy (anonymous):

how to integrate sin θ dθ/(cos^2 θ +1)

OpenStudy (mimi_x3):

\[\int\limits\frac{\sin\theta}{\cos^{2}\theta+1} d\theta \] let u = cosx =>\(\frac{du}{-sinx} \) so.. \[\int\limits\frac{-1}{u^{2}+1} \] Then integrate it..its an inverse trig function.

OpenStudy (mimi_x3):

woops typo; let u = \(\cos\theta \) => \(\frac{du}{-\sin\theta} \)

OpenStudy (dumbcow):

then you have to use the substitution u = tan du = sec^2

OpenStudy (mimi_x3):

um..why?

OpenStudy (dumbcow):

to integrate 1/(u^2 +1)

OpenStudy (anonymous):

Mimi, Sorry to ask an off topic question, but did you write those equations using Latex?

OpenStudy (mimi_x3):

Its an inverse trig function. \[\int\limits\frac{1}{a^{2}+x^{2}} dx = \frac{1}{a} \tan^{-1} \frac{x}{a} \]

OpenStudy (mimi_x3):

Yes, I did.

OpenStudy (dumbcow):

oh wait, sorry i forgot, was going to solve it from scratch :)

OpenStudy (anonymous):

Mimi, Very cool. Thanks.

OpenStudy (mimi_x3):

Oh, it will be harder i suppose..